Wednesday, May 3, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. Document on helium limits for PMTs by Dave Mack.
  2. Triggers required for the upcoming experiments
  3. Commissioning plan updates from detector systems
  4. Status update from detector owners, calibrations, efficiencies, etc.
  5. Review systems owners and update information
  6. Address system specific points of interest for the Fall 2017 run

Minutes

General

  • Dave M. showed some slides regarding helium contamination in Hall C and it's effect on PMT's installed in the Hall
  • 2.2 and 6.6 GeV is needed for detector, optics, and acceptance checkout during the Fall 2017 commissioning run
  • We have 1 week to perform the commissioning
    • There is no time to take our time with performing checkouts
    • Everything that is required to perform the checkouts needs to be available and planned out prior to commissioning
  • We need to develop a plan for measuring the trigger efficiencies
  • The trigger/DAQ configuration requirements for the upcoming experiments need to be provided as soon as possible to the trigger/DAQ group
  • Preliminary commissioning plans have been requested from each of the individual detector groups by next meeting

Beam Line

  • The super harps have been pulled from the Hall to be re-fiducialized
  • Dave G. contacted John and Kieth regarding the bug BPMs
    • They had the wrong sign during the Spring 2017 KPP run
    • This was known since work was being done on the electronics and software side
  • The known helium leak in the dump that was discovered during the KPP run is being investigated
    • The leak limits the current that can be delivered into the Hall and must be resolved prior to the Fall commissioning run
  • The leak cannot be observed now however, the helium flow has just been restored so they will check again after some time has elapsed
  • Everything else in the beam line appears to be good to go

Trigger/DAQ

  • The electronic dead time measurement is currently roughed in although it needs to be checked throughly
    • Richard Mohring's thesis was mentioned

HMS

6 GeV Drift Chambers

  • No update provided

12 GeV Drift Chambers

  • Students will become available some time in the next couple of weeks
    • Once available, work will resume

Hodoscopes

  • No update provided

Aerogel

  • No update provided

Heavy Gas Cherenkov

  • No update provided

Shower

SHMS

  • No update provided

GEM

  • No update provided

Nobel Gas Cherenkov

  • No update provided

Drift Chambers

  • No update provided

Hodoscopes

  • Simona provided a nice update, more details to follow...

Heavy Gas Cherenkov

  • No update provided

Aerogel Cerenkov

  • No update provided

Shower/Pre-shower

  • No update provided